OpenStudy (anonymous):

Factorize: (x+2)(x+3) - (x+3)(2-x)

OpenStudy (anonymous):

(x+3)(x+2)-(x+3)(2-x)= (x+3)(x+2-2+x)= (x+3)(2x)

OpenStudy (anonymous):

TY J!

OpenStudy (anonymous):

no prob